WRITE THE EQUATION OF THE LINE IN SLOPE-INTERCEPT FORM.
PASSING THROUGH THE POINTS (-1,5),(-3,-2)

Answer by joecbaseball(37)

Remember that the slope-intercept form is when the line is represented in the form:
Y = mx + b, where m is the slope and b is the y-intercept.
Now let’s figure what those values are:
First of all, your slope (m) is the ratio of the difference in the y terms divided by the difference in the x terms, or, as some teachers say, “the rise over the run”.
You have your two points, (-1,5), (-3,-2)
The difference in your y terms is (5 – (-2)) = 7
The difference in your x terms is ((-1) – (-3)) = 2
So, your slope (m) is 7/2.
Now plug that term into the generic slope-intercept form of the equation:
y = (7/2)x + b.
We now have to solve for b, and we have two different point coordinates to choose from. I’ll use (-1, 5)
Putting (-1) in for x, and 5 in for y, we get:
5 = (7/2)(-1) + b, and solve for b. This gives us that b = 17/2.
So the equation of the line for any point on the line is:
y = (7/2)x + (17/2).
Good luck